OpenStudy (firejay5):

Find the area of each triangle. Leave your answer in simplest radical form. Show work/explain answer to get medal! :D

OpenStudy (firejay5):



jhonyy9 (jhonyy9):

hope you know it sure formula for area of a triangle ,yes ?

OpenStudy (firejay5):

yes I guess I don't know what you mean

jhonyy9 (jhonyy9):

so than from this formula what date you dont know again ? so i think the height of this triangle yes ? but there you know the base what is halfed by height so than there is a right triangle in case of you need using pythagora ok ?

jhonyy9 (jhonyy9):

this is clearly now ?

OpenStudy (firejay5):

yes that's clear, but how do I leave it in simplest radical form

jhonyy9 (jhonyy9):

first of all you need calculi using pythagora the height of this triangle

OpenStudy (firejay5):

okay

OpenStudy (firejay5):

is the height 6cm

jhonyy9 (jhonyy9):

not is right 6 cm how you have got it ?

OpenStudy (firejay5):

Can you show the work of how to solve the problem or otherwise I get confused more and I learn better when work is shown

jhonyy9 (jhonyy9):

that the hypothenuse is 8 and the half part of the base is 6 so than this mean using pythagora that 8^2 = 6^2 + (height)^2 so from this how many will be the height ?

jhonyy9 (jhonyy9):

ok so from this what i ahve wrote up how many will be the height ?

OpenStudy (firejay5):

5.3cm

jhonyy9 (jhonyy9):

h^2 = 64 - 36 = 28 h=sqrt28 = 2sqrt7 than now you know the height of this triangle so now you know every dates what you need for you can calculi area of triangle

OpenStudy (firejay5):

2 sqrt 7 = 5.29 or 5.3

jhonyy9 (jhonyy9):

but you need using the radical form or not this you have wrote there up ?

OpenStudy (firejay5):

oh yea sorry